A Type II superconductor is characterized by value of London penetration depth to coherence length being :­

1) 0

2) 1

3) less than 1

4) more than 1


1
Expert's answer
2021-06-28T06:35:04-0400

Gives

Coherent length

"\xi_0=\frac{2\hbar V_f}{\pi E_g}"

Penetration depth

"\lambda _L=\sqrt{\frac{m}{\mu_0nq^2}}"

Type "(II)" superconductor

Characteristics value of London penetration depth and coherent length


"\frac{\lambda}{\xi_0}=\frac{1}{\sqrt2}=0.707"

"\frac{\lambda}{\xi_0}=0.707<1"

Option (3) is correct option